2. A particle of charge Q1 = 2μc is placed at the origin, an identical particle of charge Q2, is placed 2m from the origin on the +x-axis, and a third identical particle of charge Q3, is placed 2m from the origin on the +y-axis. 

Calculate:  

i. The net electrostatic force on Q1 ii. The net electric field on Q2 iii. The net electric potential on Q3 



1
Expert's answer
2021-07-09T08:31:23-0400

1)

"F_0=\frac {kq^2\sqrt 2}{r^2}=12.7~mN,"

2)

"E_0=\sqrt{E_1^2+E_2^2+E_1E_2\cos 45°}=\frac{kq}{2r^2}\sqrt{5+\sqrt 2}=5.7~\frac{kN}{C},"

3)

"\varphi_0=\frac{kq}r-\frac{kq}{r\sqrt 2}=2.6~kV."